OpenStudy (anonymous):

a box contains 24 cans of cola. there are 2 layers of cans and each layer holds 3 rows of 4 cans. the box is 32 cm long, 24 cm wide and 28 cm high find the height and radius of the cans find the total volume of cola in the box. answer correct to the nearest cm how many liters of cola are in the box? answer correct to 1 decimal place.

OpenStudy (anonymous):

So each can has a diameter of 8cm, based on what you've said. If the height of the box is 28cm, that means each can has a height of 14 cm. So recap: r=d/2=4cm. h=14cm. That means the volume of each = pi*r^2, which is about 703.7 cm^3. if the volume of each can is that much, than we can determine the value of all 24 (16888.7) 1 cm^3=1mL because the density of soda is about equal to that of water, therefore, 16888.7cm^3=16889mL=16.9L